### v1

@@ Introduction
The approach of modelling the banking sector as a network of interconnected entities has received increasing attention in recent years [17]. The increased attention is driven by the demand to better understand the complexity of interrelations between banks after the 2008 financial crisis, where mechanisms of an interwoven financial system were difficult to understand and predict. Consequently, more and more granular data [^2] In the given context, granular data refers to detailed, specific, and finely segmented information about individual components or entities within the financial and banking system. This data provides a detailed view of individual banks, financial institutions, or other entities, including their transactions, interactions, and various aspects of their operations. about the financial and banking system is made available to supervisors to fuel analytical models capturing interdependencies in the market.
@@
As argued by [2], the use of network theory can help in understanding financial systems, the interactions among agents and certain economic phenomena. According to [4], the complex network of exposures among banks can be considered as one of the key features for determining systemic risk while [13] review network-based approaches as analytical basis for systemic risk indicators. Recent work focuses in particular on multi-layer networks [5]. Multi-layer networks cater for the different types of interrelation among banks and model them separately in distinct layers of the network, where each layer may exhibit a different topology [6].
Empirical work analysing interbank networks is mainly driven by the availability of data. Relevant papers range from descriptive approaches to investigating the network topology to applying novel analytical frameworks to the data. [10] analysed Austrian banks based on general network theory methods and showed that the degree distribution in the network follows a powerlaw distribution, that the network has a low clustering coefficient and a short average path length. Key insights regarding the topology of an emerging network are that a core-periphery structure describes the network better than other generative models [27]. Empirical investigations of multi-layer networks capture different types of connections between banks. [5] distinguish between maturity and the secured or unsecured nature of the contracts among banks on the Italian market and showed how the resulting layers are of quite different topology. Moreover, a flattened and aggregated representation of all layers is not suitable to capture the full complexity of the network. [21] develop their model for assessing systemic risk on a multi-layer network of short-term and long-term loans and the common exposures to financial assets due to overlapping portfolios. [1] use data on exposures between large European banks, broken down by maturity and instrument type to characterise the main features of a multi-layer network.
@@
We follow [6] and define a multi-layer network as a network where each node appears in a set of different layers $\ell$, and each layer describes all the edges of a given type. In our case, the edges $e_{ij}^{\ell}$ from node $n_{i}$ to node $n_{j}$ are directed and their weight indicates the aggregated exposure of banking group $n_{i}$ towards $n_{j}$ within layer $\ell$ [^10] Note that intra-group exposures of banking groups are removed to avoid self-loops within the network.. The edge values differ across the layers $\ell$ based on the type of exposure they represent:
Long-term credit layer: This layer contains loan exposures between banking groups having an initial maturity of at least three months, in line with [21]. Exposures are aggregated over different types of loan instruments (e.g., deposits, credit lines, convenience credit, etc.), with the bulk (more than 90%) of exposures coming from deposits, credit lines other than revolving credit and unspecified loans. Edge weights $e_{ij}^{\textit{ltc}}$ indicate the outstanding nominal values.
@@
Flattened network layer: The flattened network layer is an artificially constructed layer, in the sense that it does not represent a specific real-world market. Instead, the layer is defined as the aggregation of all of the layers of the multi-layer network. More specific, we follow [6] and create a weighted aggregated overlapping adjacency matrix to form the flattened network layer.
@@
To dig deeper into the topology of the constructed multi-layer network, we leverage both graph statistics, which describe the overall network structure, and centrality measures aimed at characterizing individual nodes within the system. As seen in related work [5], this is of particular interest, as the different semantics of the layers in a multi-layer network entail different topologies and structures. The use of graph statistics jointly with centrality measures is necessary, as congruence of topology between two or more layers (graph statistics) does not necessarily imply point-wise similarity between the nodes in the layers (centrality measures) [14].
@@
Table 1 presents an overview of the global graph metrics across the different layers. The multi-layer network encompasses $114$ banking groups interacting across its various layers. Notably, the overlapping portfolio layer exhibits significant interconnectedness, evidenced by 3,614 edges. In contrast, both the short-term and long-term interbank credit layers, along with the short-term funding layer, display sparser connectivity, with densities ranging from 4% to 7%. The cross-securities layer falls between these extremes, featuring 2,456 interlinkages among the agents.141414 To have a meaningful comparison between the layers, we transform the undirected overlapping portfolio layer into a directed symmetric graph.
Table 1: Comparison of graph statistics across layers
Graph statistic Short-term credit Long-term credit Cross-securities Short-term funding Overlapping portfolio Overlapping portfolio (directed) Flattened Number of nodes 114.00 114.00 114.00 114.00 114.00 114.00 114.00 Number of edges 525.00 901.00 2,456.00 900.00 3,614.00 3,614.00 2,969.00 Number of connected components 26.00 14.00 11.00 42.00 13.00 13.00 4.00 Proportion of nodes (largest component) 0.78 0.88 0.91 0.64 0.89 0.89 0.97 Proportion of edges (largest component) 0.79 0.76 0.77 0.65 1.00 1.00 0.71 Diameter (largest component) 5.00 4.00 5.00 4.00 2.00 2.00 4.00 Avg. clustering coefficient 0.29 0.38 0.51 0.41 0.79 0.39 0.62 Reciprocity 0.42 0.48 0.47 0.71 0.00 0.00 0.58 Density 0.04 0.07 0.19 0.07 0.56 0.28 0.23 Global efficiency 0.30 0.40 0.55 0.24 0.68 0.68 0.63 Herfindahl index 0.05 0.04 0.02 0.04 0.07 0.07 0.04
@@
Turning to the clustering of nodes, Table 1 displays the average clustering coefficients across the markets. The clustering coefficient ranges from 29% (short-term credit layer) to 51% (cross-securities layer). [^16] It is worth mentioning that the clustering coefficient of the overlapping portfolio layer is substantially higher due to its undirected nature. Assuming a directed graph, the clustering coefficient aligns more closely with other layers at 39%. These values are consistent with empirical observations in other interbank markets ([24]; [5]). In conjunction with the low diameters, all layers demonstrate potential small-world properties as defined by [30]. To explore this further, we follow the approach of [12] and analyze the relationship between degree distributions and clustering coefficients across the layers. Figure 1 illustrates the results. Despite some noise, an overall negative relationship between these metrics is apparent. This suggests that highly connected agents tend to interact with counterparties who have limited interactions with each other, typically smaller entities. This observation indicates the presence of smaller-sized hubs connecting diverse market participants. In contrast, agents with lower degrees are interconnected with counterparties who are densely connected with each other, encompassing varying entity sizes, including the largest agents. This configuration hints at the existence of stable and robust relationships among certain agents. Notably, only in the overlapping portfolio layer can we confirm that the clustering coefficient is bounded away from zero. Thus, unlike other layers and in conjunction with the graph statistics, the overlapping portfolio layer stands out as a potential small-world graph.
Figure 1: Relationship between the degree ($k$) and clustering coefficient ($cc$) of banking groups within the multi-layer network. Each layer is color-coded, and the size of the markers is proportional to the total assets of the banking groups. The figure has been created using [16].
@@
Theoretical literature on interbank networks typically assume so-called scale-free networks when simulating plausible interbank networks (e.g.; [15], [22] and [23]) These types of networks commonly rely on specific statistical properties of the tails of the degree distributions underlying the networks. Technically, scale-free networks are loosely described by a probability density function which can be modelled by a power law functional relation, at least asymptotically. In this section, we examine the multi-layer network to determine whether these assumptions in literature are in line with empirical findings and how these findings might change across the different layers in the network. The premise about the potential of scale-free networks to capture the underlying network topology of interbank markets holds great importance for economic policy [21]. Contagion models often rely on scale-free topologies to simulate an interbank network (e.g. [23]), which has direct consequences on various policy implications [7].
@@
We consider the set of statistical distributions depicted in Table LABEL:tab2 as the candidates to fit $K_{weighted}^{in}(i)$ [^17] The third column of Table LABEL:tab2 shows the normalization constant $C$ such that $\sum_{x=x_{xmin}}^{\infty\ \ }{Cf(x)}=1$. The power law distribution is of main interest as this gives most evidence of the existence of scale-free networks and is usually embedded in the generating process when simulating such networks [11]. We also consider the truncated power law which behaves similar to the power law but converges towards an exponential distribution as we move deeper into the tails. Next, we consider the lognormal distribution due to its flexibility of capturing heavy tails [27]. Lastly, we consider the exponential distribution which is, per definition, the minimum alternative candidate for evaluating the existence of heavy-tails, as heavy-tailed distributions are usually defined as not exponentially bounded [19].
@@
Power law $x^{-\alpha}$ $\frac{1}{\zeta\left(\alpha,\ x_{\min}\right)}$
Truncated power law $x^{-\alpha}e^{-\lambda x}$ $\frac{\lambda^{1-\alpha}}{\Gamma\left(1-\alpha,\lambda x_{min})\right\}}$
Lognormal $\frac{1}{x}exp(-\frac{\left(\ln\left(x\right)-\ \mu\right)^{2}}{2\sigma^{2}})$ $\sqrt{\frac{2}{\left(\pi\sigma^{2}\right)}}[erfc{(\frac{\ln{\left(x_{\min}\right)-\mu}}{\sqrt{2}\sigma})]}^{-1}$
@@
Figure 2 plots the empirical probability density functions (PDFs, blue markers). For each of the different layers in the multi-layer network we fit the candidate distributions using the method of maximum likelihood. In other words, we utilize the respective Maximum Likelihood Estimators (MLEs) of the candidates to find the best fitting parameter(s) to describe the empirical observations. The left column of Figure 2 overlays the fitted candidate distributions when using all of the empirical data. The right column of Figure 2 represents the same information as the first column, but here we use the optimization scheme of [11] to fit a (truncated) power law to only the (estimated) tail of the empirical distributions.
Figure 2: The figure shows the empirical probability density functions (PDFs) over the entire data-range (left column) as well as their tail-only estimation (second column). Empirical PDFs are estimated using logarithmic binning in combination with the normalizing scheme of [20]. Empirical PDFs are depicted by the centre of these logarithmic bins as blue markers presented on logarithmic spaced axes (log-log axes). Fitted candidate distributions are presented as colour coded lines. Candidate distributions are fitted by following the methodology depicted in [11], and utilizing the Python libraries offered by [28] and [3]. The figure has been created using [16].
@@
${\mathcal{R}(p_{1},p_{2})}{=\ \sum_{i=1}^{N}{[\ln{p_{1}\left(x_{i}\right)-\ln{p_{2}\left(x_{i}\right)}}]}}.$
With $p_{1}(x)$ and $p_{2}(x)$ the PDFs of the theoretical distributions and $N$ the total number of empirical observations. A positive value indicates that $p_{1}(x)$ fits the data better than $p_{2}(x)$. To determine the best fitting distribution out of all candidates, we employ a similar methodology as [26], where we compute the following scores:
[…]
@@
Table 3 presents the results of fitting the distributions over the entire data range, encompassing both the bulk and tail parts. Across all layers, the truncated power law consistently outperforms the other three candidate distributions. Upon examination of the last column of Table 3, it is evident that while the truncated power law yields the highest fit score, the lognormal distribution also proves to be a viable candidate. Despite a lower overall score, the likelihood ratio test suggests that the lognormal distribution might offer a better fit than the truncated power law. Notably, when focusing solely on tail estimations, unreported results indicate a stronger preference for the truncated power law over the traditional power law. The added flexibility introduced by the $\lambda$ parameter enables the truncated power law to effectively account for finite-size effects near the upper edge of the in-degree exposure distribution. These findings align with those of [26], who observed that both the truncated power law and the lognormal distribution are capable of describing a wide range of interbank network measures.
Table 3: Best fitting distributions for bulk+tail estimations
Layer Best Fit Score Runner-up LR Test
Short-term Credit Truncated Power Law 27.47 Lognormal -0.53**
Long-term Credit Truncated Power Law 41.43 Lognormal -17.47***
Cross-securities Truncated Power Law 57.62 Lognormal -7.79***
Repo Market Truncated Power Law 18.93 Lognormal -2.48*
Overlapping Portfolio Truncated Power Law 129.80 Lognormal -11.75***
Flattened Network Truncated Power Law 48.81 Lognormal -21.73***
  • The last column depicts the Likelihood Ratio Test results ($\mathcal{R}$) between the winning and runner-up distributions. ***, **, and * correspond to statistical significance at the 1%, 5%, and 10% levels, respectively, based on the method introduced by [29].
@@
Centrality measures assess the centrality or degree of connection of a node within a specific layer of a multi-layer network. Typically, these metrics are derived from factors such as the number of connections, the significance of these connections, and the nodes they link, representing a node’s importance in the overall network structure [9]. While these measures may not be identical, they bear a resemblance to the concept of systemic importance extensively studied in finance and economics. Quantifying the primary drivers of systemic risk and attributing risk to individual contributors pose significant challenges [25]. Utilizing graph-based models for interbank markets and leveraging centrality measures present a promising approach to address some of these challenges. As emphasized by [17], theoretical literature on interbank networks offers a coherent method for studying interconnections, contagion processes, and systemic risk, albeit with certain limitations. Nevertheless, centrality measures can offer supplementary insights into the systemic significance of banks, surpassing traditional financial stability metrics.
@@
In general, centrality measures yield the flexibility to introduce a ranking of nodes according to their tenor within a layer, allowing one to discover the most pivotal of nodes within the system [8]. Furthermore, a comparison of a centrality measure of the same node across different layers provides insights on the differences in the topology of layers. If a node is considered very central in one layer, but less central in an another, then this underlines the different semantics of the layers.
@@
We examine the cross-correlation patterns of various centrality measures across the layers. Figure 4 shows the Kendall Rank correlation matrices. The Kendall Rank correlation is a statistic used to measure the ordinal association between two measured quantities [18]. The Kendall correlation between two evaluators will be high (with an upper bound of 1) when observations have a similar ranking between the two evaluators and low (lower bound of -1) when observations have a dissimilar ranking between the evaluators. For our cause, we appoint the different layers as the evaluators, and analyze their agreement of which nodes are most central. The first observation from Figure 4 is that strictly positive correlations are observed for all centrality measures, hinting at some agreement between the layers. On a closer look, we see that most correlation values are between 30-60%, with an exception of the ranking correlation between the cross-securities market and the flattened network, which exceed the 80% mark for most centrality measures. Another exception is the Hubs measure for which low agreement is observed between most layers.

### v2

@@ Introduction
The approach of modelling the banking sector as a network of interconnected entities has received increasing attention in recent years [huser2015too]. The increased attention is driven by the demand to better understand the complexity of interrelations between banks after the 2008 financial crisis, where mechanisms of an interwoven financial system were difficult to understand and predict. Consequently, more and more granular data [^2] In the given context, granular data refers to detailed, specific, and finely segmented information about individual components or entities within the financial and banking system. This data provides a detailed view of individual banks, financial institutions, or other entities, including their transactions, interactions, and various aspects of their operations. about the financial and banking system is made available to supervisors to fuel analytical models capturing interdependencies in the market.
@@
As argued by [allen2009networks], the use of network theory can help in understanding financial systems, the interactions among agents and certain economic phenomena. According to [bandt2012systemic], the complex network of exposures among banks can be considered as one of the key features for determining systemic risk while [di2018survey] review network-based approaches as analytical basis for systemic risk indicators. Recent work focuses in particular on multi-layer networks [bargigli2015multiplex]. Multi-layer networks cater for the different types of interrelation among banks and model them separately in distinct layers of the network, where each layer may exhibit a different topology [battiston2014structural].
Empirical work analysing interbank networks is mainly driven by the availability of data. Relevant papers range from descriptive approaches to investigating the network topology to applying novel analytical frameworks to the data. [boss2004network] analysed Austrian banks based on general network theory methods and showed that the degree distribution in the network follows a powerlaw distribution, that the network has a low clustering coefficient and a short average path length. Key insights regarding the topology of an emerging network are that a core-periphery structure describes the network better than other generative models [van2014finding]. Empirical investigations of multi-layer networks capture different types of connections between banks. [bargigli2015multiplex] distinguish between maturity and the secured or unsecured nature of the contracts among banks on the Italian market and showed how the resulting layers are of quite different topology. Moreover, a flattened and aggregated representation of all layers is not suitable to capture the full complexity of the network. [montagna2016multi] develop their model for assessing systemic risk on a multi-layer network of short-term and long-term loans and the common exposures to financial assets due to overlapping portfolios. [aldasoro2018multiplex] use data on exposures between large European banks, broken down by maturity and instrument type to characterise the main features of a multi-layer network.
@@
We follow [battiston2014structural] and define a multi-layer network as a network where each node appears in a set of different layers $\ell$, and each layer describes all the edges of a given type. In our case, the edges $e_{ij}^{\ell}$ from node $n_{i}$ to node $n_{j}$ are directed and their weight indicates the aggregated exposure of banking group $n_{i}$ towards $n_{j}$ within layer $\ell$ [^10] Note that intra-group exposures of banking groups are removed to avoid self-loops within the network.. The edge values differ across the layers $\ell$ based on the type of exposure they represent:
Long-term credit layer: This layer contains loan exposures between banking groups having an initial maturity of at least three months, in line with [montagna2016multi]. Exposures are aggregated over different types of loan instruments (e.g., deposits, credit lines, convenience credit, etc.), with the bulk (more than 90%) of exposures coming from deposits, credit lines other than revolving credit and unspecified loans. Edge weights $e_{ij}^{\textit{ltc}}$ indicate the outstanding nominal values.
@@
Flattened network layer: The flattened network layer is an artificially constructed layer, in the sense that it does not represent a specific real-world market. Instead, the layer is defined as the aggregation of all of the layers of the multi-layer network. More specific, we follow [battiston2014structural] and create a weighted aggregated overlapping adjacency matrix to form the flattened network layer.
@@
To dig deeper into the topology of the constructed multi-layer network, we leverage both graph statistics, which describe the overall network structure, and centrality measures aimed at characterizing individual nodes within the system. As seen in related work [bargigli2015multiplex], this is of particular interest, as the different semantics of the layers in a multi-layer network entail different topologies and structures. The use of graph statistics jointly with centrality measures is necessary, as congruence of topology between two or more layers (graph statistics) does not necessarily imply point-wise similarity between the nodes in the layers (centrality measures) [Engel2021].
@@
Table 1 presents an overview of the global graph metrics across the different layers. The multi-layer network encompasses $114$ banking groups interacting across its various layers. Notably, the overlapping portfolio layer exhibits significant interconnectedness, evidenced by 3,614 edges. In contrast, both the short-term and long-term interbank credit layers, along with the short-term funding layer, display sparser connectivity, with densities ranging from 4% to 7%. The cross-securities layer falls between these extremes, featuring 2,456 interlinkages among the agents. [^14] To have a meaningful comparison between the layers, we transform the undirected overlapping portfolio layer into a directed symmetric graph.
Table 1: Graph statistics across network layers
ST credit LT credit Cross-sec. ST funding Overlap Overlap (dir.) Flattened Nodes ($N$) 114 114 114 114 114 114 114 Edges ($E$) 525 901 2,456 900 3,614 3,614 2,969 Connected components 26 14 11 42 13 13 4 Largest component: share of nodes 0.78 0.88 0.91 0.64 0.89 0.89 0.97 Largest component: share of edges 0.79 0.76 0.77 0.65 1.00 1.00 0.71 Diameter (largest component) 5 4 5 4 2 2 4 Avg. clustering coefficient 0.29 0.38 0.51 0.41 0.79 0.39 0.62 Reciprocity 0.42 0.48 0.47 0.71 0.00 0.00 0.58 Density 0.04 0.07 0.19 0.07 0.56 0.28 0.23 Global efficiency 0.30 0.40 0.55 0.24 0.68 0.68 0.63 Herfindahl index 0.05 0.04 0.02 0.04 0.07 0.07 0.04
Notes: ST = short-term; LT = long-term; Cross-sec. = cross-securities; dir. = directed (symmetrized) version of the overlapping-portfolio layer.
@@
Turning to the clustering of nodes, Table 1 displays the average clustering coefficients across the markets. The clustering coefficient ranges from 29% (short-term credit layer) to 51% (cross-securities layer). [^16] It is worth mentioning that the clustering coefficient of the overlapping portfolio layer is substantially higher due to its undirected nature. Assuming a directed graph, the clustering coefficient aligns more closely with other layers at 39%. These values are consistent with empirical observations in other interbank markets ([Sor2007]; [bargigli2015multiplex]). In conjunction with the low diameters, all layers demonstrate potential small-world properties as defined by [Watts1998]. To explore this further, we follow the approach of [ContMoussaSantos] and analyze the relationship between degree distributions and clustering coefficients across the layers. Figure 1 illustrates the results. Despite some noise, an overall negative relationship between these metrics is apparent. This suggests that highly connected agents tend to interact with counterparties who have limited interactions with each other, typically smaller entities. This observation indicates the presence of smaller-sized hubs connecting diverse market participants. In contrast, agents with lower degrees are interconnected with counterparties who are densely connected with each other, encompassing varying entity sizes, including the largest agents. This configuration hints at the existence of stable and robust relationships among certain agents. Notably, only in the overlapping portfolio layer can we confirm that the clustering coefficient is bounded away from zero. Thus, unlike other layers and in conjunction with the graph statistics, the overlapping portfolio layer stands out as a potential small-world graph.
Figure 1: Relationship between the degree ($k$) and clustering coefficient (cccc) of banking groups within the multi-layer network. Each layer is color-coded, and the size of the markers is proportional to the total assets of the banking groups. The figure has been created using [Hunter:2007].
@@
Theoretical literature on interbank networks typically assume so-called scale-free networks when simulating plausible interbank networks (e.g.; [GonzlezAvella2016], [Montagna2016] and [Soramki2013]) These types of networks commonly rely on specific statistical properties of the tails of the degree distributions underlying the networks. Technically, scale-free networks are loosely described by a probability density function which can be modelled by a power law functional relation, at least asymptotically. In this section, we examine the multi-layer network to determine whether these assumptions in literature are in line with empirical findings and how these findings might change across the different layers in the network. The premise about the potential of scale-free networks to capture the underlying network topology of interbank markets holds great importance for economic policy [montagna2016multi]. Contagion models often rely on scale-free topologies to simulate an interbank network (e.g. [Soramki2013]), which has direct consequences on various policy implications [battiston2012debtrank].
@@
We consider the set of statistical distributions depicted in Table LABEL:tab2 as the candidates to fit $K_{weighted}^{in}(i)$ [^17] The third column of Table LABEL:tab2 shows the normalization constant $C$ such that $\sum_{x=x_{xmin}}^{\infty\ \ }{Cf(x)}=1$. The power law distribution is of main interest as this gives most evidence of the existence of scale-free networks and is usually embedded in the generating process when simulating such networks [Clauset2009]. We also consider the truncated power law which behaves similar to the power law but converges towards an exponential distribution as we move deeper into the tails. Next, we consider the lognormal distribution due to its flexibility of capturing heavy tails [van2014finding]. Lastly, we consider the exponential distribution which is, per definition, the minimum alternative candidate for evaluating the existence of heavy-tails, as heavy-tailed distributions are usually defined as not exponentially bounded [Markovich2018].
@@
Power law $x^{-\alpha}$ $\dfrac{1}{\zeta\!\left(\alpha,\ x_{\min}\right)}$
Truncated power law $x^{-\alpha}e^{-\lambda x}$ $\dfrac{\lambda^{1-\alpha}}{\Gamma\!\left(1-\alpha,\lambda x_{\min}\right)}$
Lognormal $\dfrac{1}{x}\exp\!\left(-\dfrac{\left(\ln x-\mu\right)^{2}}{2\sigma^{2}}\right)$ $\Bigl[\sigma\sqrt{2\pi}\cdot\tfrac{1}{2}\,\mathrm{erfc}\!\Bigl(\dfrac{\ln x_{\min}-\mu}{\sqrt{2}\sigma}\Bigr)\Bigr]^{-1}$
@@
Figure 2 plots the empirical probability density functions (PDFs, blue markers). For each of the different layers in the multi-layer network we fit the candidate distributions using the method of maximum likelihood. In other words, we utilize the respective Maximum Likelihood Estimators (MLEs) of the candidates to find the best fitting parameter(s) to describe the empirical observations. The left column of Figure 2 overlays the fitted candidate distributions when using all of the empirical data. The right column of Figure 2 represents the same information as the first column, but here we use the optimization scheme of [Clauset2009] to fit a (truncated) power law to only the (estimated) tail of the empirical distributions.
Figure 2: The figure shows the empirical probability density functions (PDFs) over the entire data-range (left column) as well as their tail-only estimation (second column). Empirical PDFs are estimated using logarithmic binning in combination with the normalizing scheme of [Milojevi2010]. Empirical PDFs are depicted by the centre of these logarithmic bins as blue markers presented on logarithmic spaced axes (log-log axes). Fitted candidate distributions are presented as colour coded lines. Candidate distributions are fitted by following the methodology depicted in [Clauset2009], and utilizing the Python libraries offered by [2020SciPy-NMeth] and [Alstott_2014]. The figure has been created using [Hunter:2007].
@@
${\mathcal{R}(p_{1},p_{2})}{=\ \sum_{i=1}^{N}{\bigl[\ln{p_{1}\left(x_{i}\right)-\ln{p_{2}\left(x_{i}\right)}}\bigr]}}.$
With $p_{1}(x)$ and $p_{2}(x)$ the PDFs of the theoretical distributions and $N$ the total number of empirical observations. A positive value indicates that $p_{1}(x)$ fits the data better than $p_{2}(x)$. To determine the best fitting distribution out of all candidates, we employ a similar methodology as [Vandermarliere2015], where we compute the following scores:
$g(p_{k})=\sum_{l\neq k}{\bigl[\mathcal{R}(p_{k},p_{l})\text{ if p-value}\leq 0.05\bigr]}$
@@
Table 3: Best-fitting distributions (bulk + tail)
Layer Best fit Score Runner-up LR test ($\mathcal{R}$)
Short-term credit Truncated power law 27.47 Lognormal -0.53∗∗
Long-term credit Truncated power law 41.43 Lognormal -17.47∗∗∗
Cross-securities Truncated power law 57.62 Lognormal -7.79∗∗∗
Repo market Truncated power law 18.93 Lognormal -2.48
Overlapping portfolio Truncated power law 129.80 Lognormal -11.75∗∗∗
Flattened network Truncated power law 48.81 Lognormal -21.73∗∗∗
Notes: The last column reports the likelihood-ratio statistic $\mathcal{R}$ comparing the best fit against the runner-up. ∗∗∗, ∗∗, and denote significance at the 1%, 5%, and 10% levels, respectively, following [Vuong1989].
Semi-parametric bootstrapped goodness-of-fit
Likelihood-ratio tests only indicate which candidate distribution provides the relatively better fit to the empirical data; they do not imply that the best-performing candidate is an adequate description of the data-generating process. In fact, for any empirical sample one can typically fit some theoretical distribution. To assess absolute fit, [Clauset2009] propose a semi-parametric bootstrap procedure for goodness-of-fit testing of heavy-tailed distributions, which also applies when the tail is fitted above an estimated cut-off value $\hat{x}_{\min}$.
The bootstrap procedure is summarized in Figure 3, and the results for the full empirical range (bulk + tail) are reported in Figure 4. Guided by the likelihood-ratio outcomes, we apply this scheme to test whether a truncated power law provides an adequate fit to the empirically observed degree distributions in each layer. The null hypothesis, $H_{0}$, is that the data are generated by a truncated power law. More precisely, $H_{0}$ states that the (positive) Kolmogorov–Smirnov distance between the empirical distribution and the fitted truncated power law can be attributed to statistical fluctuations under the truncated power-law data-generating process. We simulate $B$ bootstrap samples from the fitted model, re-estimate the parameters on each synthetic sample, and recompute the Kolmogorov–Smirnov distance. If the empirical distance is large relative to the bootstrap distribution (i.e., many bootstrap distances are smaller than the empirical one), then $H_{0}$ is unlikely and should be rejected. We use a conservative confidence level of 10% [^18] In this setting, higher confidence levels make it easier to confirm a truncated power-law data-generating process, whereas lower levels make confirmation more stringent. to decide on rejection.
As shown in Figure 4, for all layers we fail to reject $H_{0}$, and thus find evidence that the (weighted) degree distributions are consistent with a truncated power law. The empirical Kolmogorov–Smirnov distances are sufficiently small relative to their bootstrap counterparts to be plausibly explained by sampling variability, suggesting that these markets are governed by a truncated power-law data-generating mechanism. The layers are also broadly consistent in terms of parameter estimates, with $\alpha\in[1.75,2.25]$ and $\lambda\in[0.05,0.075]$, in line with values reported by [Vandermarliere2015] and [ContMoussaSantos].
Despite pronounced topological differences across layers, the weighted in-degree distributions display remarkable regularity. By contrast, when restricting attention to tail-only fitting, confirming a truncated power law becomes substantially more difficult. In unreported results we reject $H_{0}$ for the long-term credit, cross-securities, and flattened layers, indicating alternative data-generating processes; however, these tail-only conclusions are weaker given the limited number of observations in the extreme tail.
Figure 3: The bootstrapping algorithm follows the semi-parametric procedure of [Clauset2009] to assess the goodness-of-fit of heavy-tailed distributions. The approach can also be used when estimating a tail distribution with an estimated cut-off value $\hat{x}_{\min}$.
Figure 4: Semi-parametric bootstrapping results. Rows correspond to layers. The first column reports bootstrap Kolmogorov–Smirnov distances; the second column shows bootstrap lower-bound estimates $\hat{x}_{\min}$ (when performing tail-only fits); the remaining columns report bootstrap parameter estimates (e.g., $\alpha$ and $\lambda$) for the truncated power law. The vertical red dashed line marks the non-bootstrapped estimate from the empirical data, and the solid black line indicates the 90th percentile of the bootstrap distribution.
The conclusions above depend on the network characteristic under study. In particular, when examining unweighted degree distributions (both in- and out-degree), we observe substantially more heterogeneity across layers, with evidence consistent with distinct data-generating processes. Researchers and practitioners should therefore be careful when simulating networks to represent specific segments of the interbank system and should ensure that the chosen representation reflects the underlying market relationships they aim to proxy.
@@
Centrality measures assess the centrality or degree of connection of a node within a specific layer of a multi-layer network. Typically, these metrics are derived from factors such as the number of connections, the significance of these connections, and the nodes they link, representing a node’s importance in the overall network structure [Borgatti2006]. While these measures may not be identical, they bear a resemblance to the concept of systemic importance extensively studied in finance and economics. Quantifying the primary drivers of systemic risk and attributing risk to individual contributors pose significant challenges [Tarashev2009]. Utilizing graph-based models for interbank markets and leveraging centrality measures present a promising approach to address some of these challenges. As emphasized by [huser2015too], theoretical literature on interbank networks offers a coherent method for studying interconnections, contagion processes, and systemic risk, albeit with certain limitations. Nevertheless, centrality measures can offer supplementary insights into the systemic significance of banks, surpassing traditional financial stability metrics.
@@
In general, centrality measures yield the flexibility to introduce a ranking of nodes according to their tenor within a layer, allowing one to discover the most pivotal of nodes within the system [bloch2023centrality]. Furthermore, a comparison of a centrality measure of the same node across different layers provides insights on the differences in the topology of layers. If a node is considered very central in one layer, but less central in an another, then this underlines the different semantics of the layers.
@@
We examine the cross-correlation patterns of various centrality measures across the layers. Figure 5 shows the Kendall Rank correlation matrices. The Kendall Rank correlation is a statistic used to measure the ordinal association between two measured quantities [kendall1970rank]. The Kendall correlation between two evaluators will be high (with an upper bound of 1) when observations have a similar ranking between the two evaluators and low (lower bound of -1) when observations have a dissimilar ranking between the evaluators. For our cause, we appoint the different layers as the evaluators, and analyze their agreement of which nodes are most central. The first observation from Figure 5 is that strictly positive correlations are observed for all centrality measures, hinting at some agreement between the layers. On a closer look, we see that most correlation values are between 30-60%, with an exception of the ranking correlation between the cross-securities market and the flattened network, which exceed the 80% mark for most centrality measures. Another exception is the Hubs measure for which low agreement is observed between most layers.
Figure 5: Each matrix represents the set of pairwise Kendall rank correlations based on a certain centrality measure. Names of the layers are abbreviated: short-term credit layer (st_cred), long-term credit layer (lt_cred), cross-securities layer (cross_sec), short-term funding layer (st_fund), overlapping portfolio layer (ovrl_portfl) and the flattened layer (flat). The figure has been created using [Hunter:2007].
@@
Figure 6: The ranking of the 10 largest banking groups by the PageRank algorithm $C^{PR}(i)$ across different layers of the multi-layer network. Banking groups are color-coded and the size of the markers is proportional to the total assets of the banking groups. The figure has been created using [Hunter:2007].